# Crystal Chute Moves on Pipe Dreams

### v1

@@ 6.2. Weak Edelman-Greene insertion
Let $P$ be a Young tableau, let $P_{i}$ represent the $i^{\text{th}}$ row of $P$ from the bottom, and let $x$ be a letter in an ordered alphabet. The Edelman-Greene insertion of $x$ into $P$, denoted $P\leftarrow x$, is the tableau created by letting $x_{0}=x$, and applying the following algorithm, beginning with $i=0$:
@@
    If $x_{i}\geq z$ for all $z\in P_{i+1}$, then append $x_{i}$ to $P_{i+1}$ and proceed to step (3).
@@
We illustrate Edelman-Greene insertion in the following example, constructed such that each case of the algorithm occurs exactly once.
@@
In [AS, Section 4.2], the Edelman-Greene insertion algorithm is applied iteratively by inserting the letters of a given reduced word. Let $\mathbf{a}=a_{1}\cdots a_{p}\in R(w)$ for some $w\in S_{n}$. The insertion tableau for $\mathbf{a}$, denoted $P(\mathbf{a})$, is the tableau that results from Edelman-Greene inserting the letters $a_{1},\dots,a_{p}$ one by one, starting with an empty tableau. An example of an insertion tableau is given in Figure 6.6, demonstrating the letter-by-letter insertion of the word $\mathbf{a}=23124$.
pt2pt2pt3pt1pt3pt2pt1pt2pt2pt3pt1pt2pt2pt3pt4
@@
We now consider the lift operation on Young tableaux, originally defined by Assaf in [Ass2, Section 5.2]. Here, the operator produces tableaux of “key shape” in the sense of [AS, Definition 3.1], generalizing [Ass1, Definition 2.5].
@@
We now combine the Edelman-Greene insertion process with the lift operator. Given any reduced word $\mathbf{a}\in R(w)$, define the weak insertion tableau to be $\text{lift}(P(\mathbf{a}))$. Assaf refers to calculating $\text{lift}(P(\mathbf{a}))$ as the weak Edelman-Greene correspondence [Ass2]. Denote by $\operatorname{wt}(\text{lift}(P(\mathbf{a})))$ the weak composition whose $i^{\text{th}}$ coordinate equals the number of boxes in row $i$ of $\text{lift}(P(\mathbf{a}))$, enumerated from bottom to top.
@@
This final subsection is dedicated to the proof of Theorem 6.5, which then completes Theorem 2.16. We begin with several preparatory lemmas, and our final proposition identifies the insertion tableau obtained by applying Edelman-Greene insertion to any highest weight factorization.
@@
[…]
@@
We now run the Edelman-Greene insertion algorithm on a highest weight factorization.
@@
The three highest weight factorizations for $w=[21543]\in S_{5}$ are $(\;)(4)(3)(14),$ $(\;)(\;)(34)(13),$ and $(\;)(\;)(4)(134)$. Applying the Edelman-Greene algorithm to each of these factorizations produces the insertion tableaux pictured in Figure 6.8, recorded from left to right. Note that the result of the algorithm simply places the entries of block $(r^{i})$ in row $i$ of the tableau.
@@
Let $r\in RFC(w^{-1})$, and suppose that $r$ is a highest weight factorization. Our general strategy will be to show that when we run the Edelman-Greene insertion algorithm to obtain the insertion tableau $P(r)$, inserting the letter $r_{j}^{i}$ replaces $r_{j}^{i+1}$ via step (2a) by bumping all $r_{j}^{\ell}$ for $\ell\geq i+1$ up one row within the same column.
We proceed first by induction on the number of nonempty blocks in $r$. Suppose that $r=(\;)\cdots(\;)(r^{1})$ has only one nonempty block, having entries $r_{1}^{1}<\cdots<r^{1}_{m}$. Then case (1) applies in every iteration of the Edelman-Greene insertion algorithm, and the entries of block $(r^{1})$ are simply placed in row 1 of the tableau, verifying the base case.
Now suppose that $r$ has at least two nonempty blocks, and proceed to the induction step. Let $r^{\prime}=(r^{n-1})(\cdots)(r^{i+1})$, so that the insertion tableau $P(r^{\prime})$ is is obtained by placing the entries of blocks $(r^{i+1}),\dots,(r^{n-1})$ in rows $11,\dots,n-i-$ 1 of the tableau $P(r^{\prime})$, by induction on the number of blocks. We must apply the Edelman-Greene insertion algorithm using the letters of block $(r^{i})$.
@@
Let $D\in RP(w)$ be a highest weight pipe dream for $w\in S_{n}$. Denote $\Phi$ the bijection from Proposition 5.1, so that $\Phi(D)=r\in RFC(w^{-1})$ is a highest weight factorization by Corollary 5.7. By Theorem 4.10, the truncating permutation $\pi_{r}=[\pi_{1}\cdots\pi_{n}]\in S_{n}$ for the Demazure crystal $B_{\pi_{r}}(\operatorname{wt}(r))$ is the shortest permutation that sorts the weak composition $a=\operatorname{wt}(\text{lift}(P(r))$, in the sense that $(a_{\pi_{1}},\dots,a_{\pi_{n}})$ is a partition. By Proposition 6.14, however, since $r$ is a highest weight factorization, then $\operatorname{wt}(P(r))$ is a partition. Moreover, since by Proposition 6.14 the tableau $P(r)$ records the entries of block $(r^{i})$ in row $i$, then $\operatorname{wt}(P(r))=\operatorname{wt}(r)$. Therefore, Theorem 4.10 says that $\operatorname{wt}(r)=(a_{\pi_{1}},\dots,a_{\pi_{n}})$, or equivalently,

### v2

@@ 6.2. Weak Edelman–Greene insertion
Let $P$ be a Young tableau, let $P_{i}$ represent the $i^{\text{th}}$ row of $P$ from the bottom, and let $x$ be a letter in an ordered alphabet. The Edelman–Greene insertion of $x$ into $P$, denoted $P\leftarrow x$, is the tableau created by letting $x_{0}=x$, and applying the following algorithm, beginning with $i=0$:
@@
    If $x_{i}\geq z$ for all $z\in P_{i+1}$, then append $x_{i}$ to $P_{i+1}$, and the algorithm terminates.
@@
We illustrate Edelman–Greene insertion in the following example, constructed such that each case of the algorithm occurs exactly once.
@@
In [AS, Section 4.2], the Edelman–Greene insertion algorithm is applied iteratively by inserting the letters of a given reduced word. Let $\mathbf{a}=a_{1}\cdots a_{p}\in R(w)$ for some $w\in S_{n}$. The insertion tableau for $\mathbf{a}$, denoted $P(\mathbf{a})$, is the tableau that results from Edelman–Greene inserting the letters $a_{1},\dots,a_{p}$ one by one, starting with an empty tableau. An example of an insertion tableau is given in Figure 6.6, demonstrating the letter-by-letter insertion of the word $\mathbf{a}=23124$.
223132122312234
@@
We now consider the lift operation on Young tableaux, originally defined by Assaf in [Ass1, Section 5.2]. Here, the operator produces tableaux of “key shape” in the sense of [AS, Definition 3.1], generalizing [Ass2, Definition 2.5].
@@
We now combine the Edelman–Greene insertion process with the lift operator. Given any reduced word $\mathbf{a}\in R(w)$, define the weak insertion tableau to be $\text{lift}(P(\mathbf{a}))$. Assaf refers to calculating $\text{lift}(P(\mathbf{a}))$ as the weak Edelman–Greene correspondence [Ass1]. Denote by $\operatorname{wt}(\text{lift}(P(\mathbf{a})))$ the weak composition whose $i^{\text{th}}$ coordinate equals the number of boxes in row $i$ of $\text{lift}(P(\mathbf{a}))$, enumerated from bottom to top.
@@
This final subsection is dedicated to the proof of Theorem 6.5, which then completes Theorem 2.16. We begin with several preparatory lemmas, and our final proposition identifies the insertion tableau obtained by applying Edelman–Greene insertion to any highest weight factorization.
@@
Let $r\in RFC(w)$ for $w\in S_{n}$ be a highest weight factorization. If $r^{i}_{j}$ denotes the $j^{\text{th}}$ element in block $(r^{i})$, then $r_{j}^{n-1}>\cdots>r_{j}^{2}>r_{j}^{1}$, where we omit letters from any empty blocks.
@@
We now run the Edelman–Greene insertion algorithm on a highest weight factorization.
@@
The three highest weight factorizations for $w=[21543]\in S_{5}$ are $(\;)(4)(3)(14),$ $(\;)(\;)(34)(13),$ and $(\;)(\;)(4)(134)$. Applying the Edelman–Greene algorithm to each of these factorizations produces the insertion tableaux pictured in Figure 6.8, recorded from left to right. Note that the result of the algorithm simply places the entries of block $(r^{i})$ in row $i$ of the tableau.
@@
Let $r\in RFC(w^{-1})$, and suppose that $r$ is a highest weight factorization. Our general strategy will be to show that when we run the Edelman–Greene insertion algorithm to obtain the insertion tableau $P(r)$, inserting the letter $r_{j}^{i}$ replaces $r_{j}^{i+1}$ via step (2a) by bumping all $r_{j}^{\ell}$ for $\ell\geq i+1$ up one row within the same column.
We proceed first by induction on the number of nonempty blocks in $r$. Suppose that $r=(\;)\cdots(\;)(r^{1})$ has only one nonempty block, having entries $r_{1}^{1}<\cdots<r^{1}_{m}$. Then case (1) applies in every iteration of the Edelman–Greene insertion algorithm, and the entries of block $(r^{1})$ are simply placed in row 1 of the tableau, verifying the base case.
Now suppose that $r$ has at least two nonempty blocks, and proceed to the induction step. Let $r^{\prime}=(r^{n-1})(\cdots)(r^{i+1})$, so that the insertion tableau $P(r^{\prime})$ is obtained by placing the entries of blocks $(r^{i+1}),\dots,(r^{n-1})$ in rows $11,\dots,n-i-$ 1 of the tableau $P(r^{\prime})$, by induction on the number of blocks. We must apply the Edelman–Greene insertion algorithm using the letters of block $(r^{i})$.
@@
Let $D\in RP(w)$ be a highest weight pipe dream for $w\in S_{n}$. Denote by $\Phi$ the bijection from Proposition 5.1, so that $\Phi(D)=r\in RFC(w^{-1})$ is a highest weight factorization by Corollary 5.7. By Theorem 4.10, the truncating permutation $\pi_{r}=[\pi_{1}\cdots\pi_{n}]\in S_{n}$ for the Demazure crystal $B_{\pi_{r}}(\operatorname{wt}(r))$ is the shortest permutation that sorts the weak composition $a=\operatorname{wt}(\text{lift}(P(r))$, in the sense that $(a_{\pi_{1}},\dots,a_{\pi_{n}})$ is a partition. By Proposition 6.14, however, since $r$ is a highest weight factorization, then $\operatorname{wt}(P(r))$ is a partition. Moreover, since by Proposition 6.14 the tableau $P(r)$ records the entries of block $(r^{i})$ in row $i$, then $\operatorname{wt}(P(r))=\operatorname{wt}(r)$. Therefore, Theorem 4.10 says that $\operatorname{wt}(r)=(a_{\pi_{1}},\dots,a_{\pi_{n}})$, or equivalently,